write the expression in a natural logarithm.
3 ln 3 + 3lnc.
I believe we need to add the like terms?

Answer by Alan3354(69443)


3 ln 3 + 3lnc.
-------------
3ln(3) + 3ln(c) = ln(3^3) + ln(c^3)
= ln(27c^3)